# ASIDE: Architectural Separation of Instructions and Data in Language Models

### v3

@@ Abstract
Despite their remarkable performance, large language models lack elementary safety features, making them susceptible to numerous malicious attacks. In particular, previous work has identified the absence of an intrinsic separation between instructions and data as a root cause of the success of prompt injection attacks. In this work, we propose a new architectural element, ASIDE, that allows language models to clearly separate instructions and data at the level of embeddings. ASIDE applies an orthogonal rotation to the embeddings of data tokens, thus creating clearly distinct representations of instructions and data tokens without introducing any additional parameters. As we demonstrate experimentally across a range of models, instruction-tuning LLMs with ASIDE (1) leads to highly increased instruction-data separation without a loss in model utility and (2) makes the models more robust to prompt injection benchmarks, even without dedicated safety training. Additionally, we provide insights into the mechanism underlying our method through an analysis of the model representations. The source code and training scripts are openly accessible at https://github.com/egozverev/aside.
@@ 1 Introduction
Large language models (LLMs) are commonly associated with interactive, open-ended chat applications, such as ChatGPT. However, in many practical applications LLMs are integrated as a component into larger software systems. Their rich natural language understanding abilities allow them to be used for text analysis and generation, translation, summarization, or information retrieval (Zhao et al., 2023). In all of these scenarios, the system is given instructions, for example as a system prompt, and data, for example, a user input or an uploaded document. These two forms of input play different roles: the instruction should be executed, determining the behavior of the model, while the data should be processed, i.e., transformed to become the output of the system. In other words, the instructions are meant to determine and maintain the function implemented by the model, whereas the data becomes the input to this function.
@@
Current LLM architectures lack a built-in mechanism that would distinguish which part of their input constitutes instructions, and which part constitutes data. Instead, the two roles are generally distinguished indirectly, e.g., by natural language statements within the prompt, or by special tokens. It is widely observed that this form of instruction-data separation is insufficient, contributing to the models’ vulnerability to many attack patterns, such as indirect prompt injection (Greshake et al., 2023) or system message extraction (Zhang et al., 2024b). As a result, current LLMs are problematic for safety-critical tasks (Anwar et al., 2024).
@@
While initial works on instruction-data separation were qualitative or exploratory in nature, Zverev et al. (2025) recently introduced a quantitative study of this phenomenon. Their experiments confirmed that none of models they tested provided a reliable separation between instructions and data, and that straightforward mitigation strategies, such as prompt engineering (Hines et al., 2024), prompt optimization (Zhou et al., 2024) or fine-tuning (Piet et al., 2024), are insufficient to solve the problem.
@@
In this work, we go one step further and not only describe the problem but offer a path towards a principled solution. We propose a new architectural element, ASIDE (Architecturally Separated Instruction-Data Embeddings), that enforces the separation between instructions and data on the level of model architecture rather than just on the level of input prompts or model weights. Our core hypothesis is that in order to achieve instruction-data separation, the model should have an explicit representation from the first layer on, which of the input tokens are executable and which are not. To achieve this, ASIDE assigns each input token one of two embedding representations based on its functional role (instruction or data). ASIDE can be integrated into existing language models without a need for repeating their pretraining. One only modifies the forward pass of the model to take the functional role of each token as an input and apply a fixed orthogonal rotation to the embeddings of data tokens. Then instruction-tuning in a standard supervised fine-tuning setup is applied.
@@
As we show experimentally, this seemingly minor change of architecture has a number of beneficial consequences. First, it allows the model to reliably determine a token’s functional role already from the first layer. This is in contrast to conventional models, which only have one embedding per token. For them, each time a token occurs, it is represented by the same embedding vector, so the token representation itself does not contain any information about its functional role. Instead, the model has to infer from its context whether a token should be executed or processed, and it must learn the ability to do so during the training (typically instruction tuning).
@@
Second, we demonstrate that even when trained on standard instruction-tuning data without dedicated safety-tuning, ASIDE models achieve better separation scores in the sense of Zverev et al. (2025) while preserving the model’s utility, as well as achieving higher robustness against prompt injection. This effect holds consistently across a variety of models, including Llama 3.1, Llama 2, Qwen 2.5, and Mistral models. Besides quantitative results, we also provide qualitative insights into ASIDE’s inner working mechanism by analyzing the models’ ability to distinguish between instruction and data representations.
@@ 2 Related work
Large language models (LLMs) face a range of vulnerabilities, including prompt injection (Chen et al., 2025a; Yi et al., 2025; Hines et al., 2024; Chen et al., 2024), goal hijacking (Perez & Ribeiro, 2022; Chen & Yao, 2024; Levi & Neumann, 2024), prompt stealing (Perez & Ribeiro, 2022; Hui et al., 2024; Yang et al., 2024), or data leakage (Carlini et al., 2021; Huang et al., 2022). See, for example, Das et al. (2024) or Yao et al. (2024) for recent surveys. Like us, Zverev et al. (2025) argue that a crucial factor towards such vulnerabilities is the lack of instruction-data separation in current models. Wallace et al. (2024) put forward the idea of an instruction hierarchy that would give some inputs a higher priority for being executed than others (with pure data located at the lowest level of the hierarchy). Existing defenses include (1) prompt engineering (Zhang et al., 2025; Hines et al., 2024; Chen & Yao, 2024; Perez & Ribeiro, 2022), (2) optimization-based techniques, such as adversarial training (Chen et al., 2025a; Piet et al., 2024; Chen et al., 2024) and circuit-breaking (Zou et al., 2024), (3) and injection detection (Microsoft, 2024; Abdelnabi et al., 2025; Chen et al., 2025b). In concurrent work, Debenedetti et al. (2025) proposed to create a protective security system to control the instruction/data flow of LLMs. This, however, operates on a system’s level outside of the LLM and is therefore orthogonal to our approach of improving instruction-data separation with the LLM itself.
At the same time, architectural solutions remain largely absent for instruction-data separation, despite their success in other areas. Li & Liang (2021) use prefix tuning to prepend task-specific embeddings that steer model behavior without altering core weights. Su et al. (2024) apply rotations to encode token positions, showing that geometric transformations can efficiently inject structural information into embeddings. These examples illustrate how embedding-level changes — especially rotations — can assist in separating functional roles of tokens. Yet applications of such techniques to safety remain unexplored. ASIDE addresses this gap by applying a fixed orthogonal rotation to data token embeddings, extending rotation-based methods to the safety domain without adding parameters or sacrificing performance.
@@ 3 Architecturally Separated Instruction-Data Embeddings
We now introduce our main contribution, the ASIDE (Architecturally Separated Instruction-Data Embeddings) method of data encoding for large language models. At the core of ASIDE lies the idea that instructions and data should have different representations. A natural place to enforce this in a language model is at the level of token embeddings: if a token’s functional role (instruction or data) can be read off from its embeddings, the model can easily maintain this distinction in the later layers’ representations. However, simply learning different embeddings for data and instruction tokens would be impractical: it would double the number of learnable parameters in the embedding layer, and training them would require a lot of (pre-)training data with annotated functional roles for all tokens, which standard web-scaped sources do not possess.
@@
In the rest of the section, we first provide the technical definition of ASIDE’s architectural component. Afterwards, we describe our suggested way for converting existing models to benefit from ASIDE without having to retrain them from scratch. Note that we target the setting in which for every token the information, which of the two cases it is, is available at input time, e.g., because instructions and data originate from different input sources, or because instructions are marked by specific tags. Alternative setups, such as when the functional role of tokens (instruction or data) is inferred at runtime, are clearly interesting and relevant, but beyond the scope of this work.
Architectural Element
@@
The main architectural component of ASIDE is a conditional embedding mechanism that takes the functional role of an input token into account. If a token is executable, i.e., part of an instruction, it is represented by a different embedding vector than if it is not executable, i.e., part of passive data. To implement the conditional embedding mechanism standard language model components suffice: let denote a model’s token embedding matrix, where is the vocabulary size and is the embedding dimensionality. For a token, , let be its index in the vocabulary. Now, ASIDE works as follows: if a token is part of the instructions, it is embedded as , as it would be in a standard architecture. However, if the same token appears as data, we apply a fixed (i.e., not learnable) orthogonal rotation to that embedding during the forward pass, resulting in an embedding . While in principle, any rotation matrix could be used, in practice we rely on an isoclinic one, which is easy to implement and efficient to perform. Specifically, the embedding dimensions are split into groups of size 2 and each of these is multiplied by a -rotation matrix .
Training
@@
Because ASIDE only modifies the embedding layer’s forward pass (rather than the embeddings themselves), it can also be integrated post hoc into any pretrained LLM. To do so, we suggest a two-step procedure: 1) modify the model’s forward pass to include the additional rotation for data tokens, 2) fine-tune the resulting model on a dataset that allows the network to learn the different roles of tokens in executable versus non-executable context.
@@
The ASIDE construction is agnostic to the underlying model architecture in the sense that it is applicable to any model that starts by a token-embedding step and it is not restricted to any specific choice of tokenizer. Furthermore, it readily allows for domain-specific extensions, such as scenarios where only a subset of tokens are role-distinguished (i.e., only certain “critical” tokens are rotated). If more than two functional levels are needed (e.g., a multi-tier instruction hierarchy), these could also be implemented by defining additional orthogonal transformations. However, we leave such extensions to future work.
@@ 4 Experiments: Instruction-data separation
Our first experimental evaluation of ASIDE models (i.e., with conditional token embeddings) studies their ability to separate instructions and data in a general instruction-following setting. We describe our training procedure in Section 4.1, the evaluation pipeline in Section 4.2, and we discuss the results in Section 4.3.
@@ 4.1 Training procedure
Models
We use Qwen 2.5 7B (Yang et al., 2025), Mistral 7B v0.3 (Jiang et al., 2023) and several generations of the Llama models (Touvron et al., 2023; Grattafiori et al., 2024): Llama 3.1 8B, Llama 2 7B, and Llama 2 13B. In all cases, we compare three architectures: the standard model without modifications (called Vanilla), the model with ASIDE modification, and the model architecture suggested in Wu et al. (2024) (called ISE). To make our baselines competitive, we use specialized tokens to mark the beginning and the end of instruction and data blocks in the input similar to Chen et al. (2024) and include a system prompt similar to the one used by Taori et al. (2023) that specifies which parts of the input are instructions and which are data. Note that we do not use models that were explicitly instruction- or safety-tuned in our experiments, but instead start from a plain pretrained model to avoid biasing the safety evaluations.
Data
@@
Model training
All models and architectures are trained using the same supervised fine-tuning procedure. The models are trained for 3 epochs. The hyperparameters (learning rate ; batch size for 7B/8B models, for 13B/14B models; warm-up ratio ) are chosen as the ones with the lowest validation loss across all runs. See Appendix B for details.
@@ 4.2 Evaluation procedure
Instruction-data separation (SEP) score
As our main quantity of interest, for each model we compute its instruction-data separation score, following the protocol of Zverev et al. (2025). We rely on the SEP dataset222https://github.com/egozverev/Should-It-Be-Executed-Or-Processed, which consists of 9160 pairs of instructions and inputs. To compute the separation score, one first takes a set of (instruction, data) pairs. Then for each pair, one puts an unrelated instruction (called probe) in either “data” or “instruction” part of the input and compares the outputs. Models achieve a high score if they execute the probes in the “instruction” part, but do not execute them in the “data” part.
Utility evaluation
We use two benchmarks for evaluating utility: the SEP Utility metric from Zverev et al. (2025), and the commonly used AlpacaEval 1.0 (Dubois et al., 2024a, b). SEP Utility measures how often the model executes instructions in the SEP dataset. AlpacaEval 1.0 employs an LLM judge (GPT-4) to measure how often the outputs of the evaluated model are preferable to GPT-3.5 (text-davinci-003).
@@ 4.3 Results
We report the main results of our evaluation in Figure 2 (and Table 3 in Appendix). In addition to the three instruction-tuned setups (Vanilla, ISE, ASIDE), we also include results for the corresponding Base models, which were pre-trained but not instruction-tuned. In all cases, ASIDE achieves significantly higher separation scores than the other methods, while achieving comparable utility values.
Specifically, we observe that ASIDE increases the SEP scores between 12.3 (Llama 2 7B) and 44.1 (Mistral 7B) percentage points (p.p.) compared to the standard (Vanilla) model. The utility values of ASIDE-models show only minor differences to the Vanilla ones, both in terms of SEP Utility as well as AlpacaEval. A single exception is SEP Utility for the Mistral 7B model, where the SEP Utility drops from 45.0% to 25.3%. This, however, we believe to be an artifact of the rather brittle utility evaluation, as ASIDE’s AlpacaEval score is slightly higher than Vanilla’s, and highest SEP Utility score in this setting is actually achieved by the non-instruction-tuned Base model.
@@
Note that in contrast to prior work, our fine-tuning procedure does not contain specific measures to increase separation or safety, neither in the optimization objective nor in the dataset. Consequently, we conclude that the increase in instruction-data separation is truly the result of the change in model architecture.

### v4

@@ Abstract
Despite their remarkable performance, large language models lack elementary safety features, making them susceptible to numerous malicious attacks. In particular, previous work has identified the absence of an intrinsic separation between instructions and data as the root cause of the success of prompt injection attacks. In this work, we propose a new architectural element, ASIDE, that allows language models to clearly separate instructions and data at the level of token embeddings. ASIDE applies an orthogonal rotation to the embeddings of data tokens, thus creating clearly distinct representations of instructions and data tokens without introducing any additional parameters. As we demonstrate experimentally across a range of models, instruction-tuning LLMs with ASIDE (1) achieves substantially higher instruction-data separation without performance loss and (2) makes the models more robust to prompt injection benchmarks, even without dedicated safety training. Additionally, we provide insights into the mechanism underlying our method through an analysis of the model representations.
@@ 1 Introduction
Large language models (LLMs) are commonly associated with interactive chat applications, such as ChatGPT. However, in many practical applications, LLMs are integrated as parts of larger software systems (Weber, 2024), such as email clients (Abdelnabi et al., 2025b) and agentic pipelines (Costa et al., 2025). Their rich natural language understanding abilities allow them to be used for text analysis and generation, translation, summarization, or information retrieval (Zhao et al., 2023).
In many of these scenarios, the system is given instructions, for example as a system prompt, and data, for example, a user input or an uploaded document. These two forms of input play different roles: the instruction should be executed, determining the behavior of the model, while the data should be processed, i.e., transformed to become the output of the system. In other words, the instructions are meant to determine and maintain the function implemented by the model, whereas the data becomes the input to this function.
@@
In other areas of computer science, the separation between executable and non-executable memory regions lies at the core of safety measures that prevent, e.g., SQL injections (Clarke-Salt, 2009) or buffer overflow exploits (Paulson, 2004). In contrast, current LLM architectures lack a built-in mechanism that would distinguish which part of their input constitutes instructions, and which part constitutes data. Instead, the two roles are generally distinguished indirectly, e.g., by natural language statements within the prompt or by special tokens (Hines et al., 2024). It is widely observed that this form of instruction-data separation is insufficient (Zverev et al., 2025), contributing to the models’ vulnerability to many attack patterns, such as indirect prompt injection (Greshake et al., 2023) or system message extraction (Zhang et al., 2024b). As a result, current LLMs are problematic for safety-critical tasks (Anwar et al., 2024).
@@
While initial works on instruction-data separation were qualitative or exploratory in nature, Zverev et al. (2025) recently presented a quantitative study of the phenomenon. Their experiments confirmed that none of the models they tested provided a reliable separation between instructions and data, and that straightforward mitigation strategies, such as prompt engineering (Hines et al., 2024), prompt optimization (Zhou et al., 2024) or fine-tuning (Piet et al., 2024), are insufficient to solve the problem.
@@
In this work, we go one step further: we not only describe the problem but offer a path towards a principled solution. We propose a new architectural element, ASIDE (Architecturally Separated Instruction-Data Embeddings), that enforces the separation between instructions and data at the level of model architecture rather than just at the level of input prompts or model weights. Our core hypothesis is that in order to achieve instruction-data separation, the model should have an explicit representation from the first layer onward, which of the input tokens are executable and which are not. To achieve this, ASIDE assigns each input token one of two embedding representations based on its functional role (instruction or data). See Figure 1 for an illustration. ASIDE can be integrated into existing language models without a need for repeating their pretraining. Only the model’s forward pass needs to be modified to accept each token’s functional role as input and to apply a fixed orthogonal rotation to data token embeddings. Then instruction-tuning in a standard supervised fine-tuning setup is applied.
@@
As we show experimentally, this seemingly minor change in the architecture has two major advantages. First, it allows the model to reliably determine a token’s functional role already from the first layer. This is in contrast to conventional models, which only have one embedding per token. For them, each time a token occurs, it is represented by the same embedding vector. The token representation itself does not contain any information about its functional role. A conventional model has to infer from the context whether a token should be executed or processed, and it must learn to do so during training.
@@
Second, even when trained on standard instruction-tuning data without dedicated safety-tuning, ASIDE models achieve better separation scores in the sense of Zverev et al. (2025) while preserving the model’s utility, as well as achieving higher robustness against prompt injection. This is achieved without adversarial training examples. This effect holds consistently across a variety of models, including Qwen 3, Qwen 2.5, Llama 3.1, Llama 2, and Mistral models. Besides quantitative results, we also provide qualitative insights into ASIDE’s inner working mechanism by analyzing the models’ ability to distinguish between instruction and data representations.
@@ 2 Related Work
Large language models (LLMs) face a range of vulnerabilities, including prompt injection (Chen et al., 2025a; Yi et al., 2025; Hines et al., 2024; Chen et al., 2024), goal hijacking (Perez and Ribeiro, 2022; Chen and Yao, 2024; Levi and Neumann, 2024), prompt stealing (Perez and Ribeiro, 2022; Hui et al., 2024; Yang et al., 2024), or data leakage (Carlini et al., 2021; Huang et al., 2022). See, for example, Das et al. (2024) or Yao et al. (2024) for recent surveys. Like us, Zverev et al. (2025) argue that a crucial factor contributing to such vulnerabilities is the lack of instruction-data separation in current models. Wallace et al. (2024) put forward the idea of an instruction hierarchy that would give some inputs a higher priority for being executed than others (with pure data located at the lowest level of the hierarchy). Existing defenses include (1) prompt engineering (Zhang et al., 2025; Hines et al., 2024; Chen and Yao, 2024; Perez and Ribeiro, 2022), (2) optimization-based techniques, such as adversarial training (Chen et al., 2025a; Piet et al., 2024; Chen et al., 2024) and circuit-breaking (Zou et al., 2024), and (3) injection detection (Microsoft, 2024; Abdelnabi et al., 2025a; Chen et al., 2025b). In concurrent work, Debenedetti et al. (2025) and Costa et al. (2025) proposed to create a protective security system to control the instruction/data flow of LLMs. This, however, operates at the system level outside of the LLM and is therefore orthogonal to our approach of improving instruction-data separation with the LLM itself.
Architectural solutions remain largely absent for instruction-data separation, despite their success in other areas. Li and Liang (2021) use prefix tuning to prepend task-specific embeddings that steer model behavior without altering core weights. Su et al. (2024) apply rotations to encode token positions, showing that geometric transformations can inject structural information into embeddings. These examples illustrate how embedding-level changes - especially rotations - can assist in separating functional roles of tokens. Yet applications of such techniques to safety remain unexplored. ASIDE addresses this gap by applying a fixed orthogonal rotation to data token embeddings, extending rotation-based methods to the safety domain without adding parameters or sacrificing performance.
@@ 3 Architecturally Separated Instruction-Data Embeddings
We now introduce our main contribution, the ASIDE (Architecturally Separated Instruction-Data Embeddings) method of data encoding for large language models. At the core of ASIDE lies the idea that instructions and data should have different representations. A natural place to enforce this in a language model is at the level of token embeddings: if a token’s functional role (instruction or data) can be read off from its embeddings, the model can easily maintain this distinction in the later layers’ representations. However, simply learning different embeddings for data and instruction tokens would be impractical: it would double the number of learnable parameters in the embedding layer, and training them would require a lot of (pre-)training data with annotated functional roles for all tokens, which standard web-scraped sources do not possess.
@@
In the rest of the section, we first provide the technical definition of ASIDE’s architectural component. Afterwards, we describe our suggested way of converting existing models to benefit from ASIDE without having to retrain them from scratch. Note that we target the setting in which the information about which of the two roles a token has is available at input time, e.g., because instructions and data originate from different input sources. This is a common situation when LLMs are used as components of larger software solutions, e.g., in an email client, where the contents of the emails should always be treated as data, not as instructions (Abdelnabi et al., 2025b). Alternative setups, for example, inferring the functional role of tokens (instruction or data) at runtime, to use in a general-purpose assistant chatbot, are interesting and relevant, but lie beyond the scope of this work.
Architectural Element.
@@
The main architectural component of ASIDE is a conditional embedding mechanism that takes the functional role of an input token into account. If a token is executable, i.e., part of an instruction, it is represented by a different embedding vector than if it is not executable, i.e., part of passive data. To implement the conditional embedding mechanism, standard language model components suffice: let denote a model’s token embedding matrix, where is the vocabulary size and is the embedding dimensionality. For a token, , let be its index in the vocabulary. Now, ASIDE works as follows: if a token is part of the instructions, it is embedded as , as it would be in a standard architecture. However, if the same token appears as data, we apply a fixed (i.e., not learnable) orthogonal rotation to that embedding during the forward pass, resulting in an embedding . While in principle, any rotation matrix could be used, in practice we rely on an isoclinic one, which is easy to implement and efficient to perform. Specifically, the embedding dimensions are split into groups of size 2 and each of these is multiplied by a -rotation matrix . See details in Appendix A.
Implementation.
@@
Because ASIDE only modifies the embedding layer’s forward pass (rather than the embeddings themselves), it can also be integrated post hoc into any pretrained LLM. To do so, we suggest a two-step procedure: 1) modify the model’s forward pass to include the additional rotation for data tokens, 2) fine-tune the resulting model on a dataset that allows the network to learn the different roles of tokens in executable versus non-executable context. We assume that token roles are fixed by system design (e.g., external files are always labeled as data). Unlike prompt-based methods (Hines et al., 2024), this prevents role hijacking via delimiters in external content. See Appendix B for details.
@@
The ASIDE construction is agnostic to the underlying model architecture in the sense that it is applicable to any model that starts with a token-embedding step and it is not restricted to any specific choice of tokenizer. Furthermore, it readily allows for domain-specific extensions, such as scenarios where only a subset of tokens are role-distinguished (i.e., only certain “critical” tokens are rotated). If more than two functional levels are needed (e.g., a multi-tier instruction hierarchy), these could also be implemented by defining additional orthogonal transformations. However, we leave such extensions to future work.
@@ 4 Experiments: Instruction-data separation
Our first experimental evaluation of ASIDE models (i.e., with conditional token embeddings) studies their ability to separate instructions and data in a general instruction-following setting.
@@ 4.1 Training procedure
Models.
We use Qwen 3 8B (Yang et al., 2025a), Qwen 2.5 7B (Yang et al., 2025b), Mistral 7B v0.3 (Jiang et al., 2023) and several generations of the Llama models (Touvron et al., 2023; Grattafiori et al., 2024): Llama 3.1 8B, Llama 2 7B, and Llama 2 13B. In all cases, we compare three model architectures:
Vanilla Architecture: Naively fine-tuning a standard architecture does not allow enforcing any separation. To make the comparisons meaningful, we therefore implement some changes during training and inference: 1) we introduce specialized tokens to mark the beginning and end of instruction and data blocks in the input, similar to Chen et al. (2024). 2) we include a prompt (similar to the one used by Taori et al. (2023)) that specifies which parts of the input are instructions and which are data.
ISE: The model architecture from Wu et al. (2024), where data embeddings are offset from instruction embeddings by a learnable vector.
ASIDE: Our proposed modification that applies an orthogonal rotation to data embeddings.
Note that we use plain pretrained models rather than instruction- or safety-tuned models to avoid biasing the safety evaluations.
Data.
@@
Model training.
All models and architectures are trained using the same supervised fine-tuning procedure. The models are trained for 3 epochs. The hyperparameters (learning rate ; batch size for 7B/8B models, for 13B/14B models; warm-up ratio ) are chosen as the ones with the lowest validation loss across all runs. See Appendix C for details.
@@ 4.2 Evaluation procedure
Instruction-data separation (SEP) score.
As our main quantity of interest, for each model we compute its instruction-data separation score, following the protocol of Zverev et al. (2025). We rely on the SEP dataset222https://github.com/egozverev/Should-It-Be-Executed-Or-Processed, which consists of 9160 pairs of instructions and inputs. To compute the separation score, one first takes a set of (instruction, data) pairs. Then for each pair, one puts an unrelated instruction (called probe) in either the “data” or the “instruction” part of the input and compares the outputs. Models achieve a high score if they execute the probes in the “instruction” part, but do not execute them in the “data” part.
Utility evaluation.
We use two benchmarks for evaluating utility: the SEP Utility metric from Zverev et al. (2025), and AlpacaEval 1.0 (Dubois et al., 2024a; b). SEP Utility measures how often the model executes instructions in the SEP dataset. AlpacaEval 1.0 employs an LLM judge (GPT-4) to measure how often the outputs of the evaluated model are preferable to GPT-3.5 (text-davinci-003).
@@ 4.3 Results
We report the results of our evaluation in Figure 2 (and Table 3 in Appendix). In addition to the three instruction-tuned setups (Vanilla, ISE, ASIDE), we also include results for the corresponding Base models, which were pre-trained but not instruction-tuned. In all cases, ASIDE achieves significantly higher separation scores than the other methods, while achieving comparable utility values.
Specifically, we observe that ASIDE increases the SEP scores between 12.3 (Llama 2 7B) and 44.1 (Mistral 7B) percentage points (p.p.) compared to the standard (Vanilla) model. The utility values of ASIDE-models show only minor differences to the Vanilla ones, both in terms of SEP Utility as well as AlpacaEval. A single exception is Mistral-7B, where SEP Utility decreases while AlpacaEval improves slightly. We believe this, however, to be an artifact of the rather brittle utility evaluation, as ASIDE’s AlpacaEval score is slightly higher than Vanilla’s, and highest SEP Utility score in this setting is actually achieved by the non-instruction-tuned Base model.
@@
Note that in contrast to prior work, our fine-tuning procedure does not contain specific measures to increase separation or safety, either in the optimization objective or in the dataset. Thus, we conclude that the increase in instruction-data separation is truly the result of the change in model architecture.